y=5x -126 when x is 3 times as great as y What is the value of x

Answer by nerdybill(7384)

y=5x -126 when x is 3 times as great as y What is the value of x
.
From:"x is 3 times as great as y" we have:
x = 3y
solving for y we get:
x/3 = y
.
substitute the above into:
y=5x -126
x/3 = 5x -126
multiplying both sides by 3:
x = 15x -378
-14x = -378
x = -378/(-14)
x = 27